Suppose that you are an investor based in Switzerland, and you expect the U.S. dollar to depreciate by 2.75 percent over the next year. The interest rate on one-year risk-free bonds is 5.25 percent in the United States and 2.75 percent in Switzerland. The current exchange rate is SFr1.62 per U.S. dollar. te the foreign currency risk premium from the Swiss investors viewpoint. that the Swiss investors expectations are met. b. Calculate the return on the U.S bond from the Swiss investors viewpoint, assuming
Answer #1

a) S0 = 1.62 SFr

E(S1) = [1.62 - (1.62*2.75%)] SFr

E(S1) = (1.62 - 0.04455) SFr

E(S1) = 1.57545 SFr

Expected appreciation in foreign currency = (1.57545 - 1.62) / 1.62

Expected appreciation in foreign currency = - 0.0275

foreign currency is going to depreciate in future

Interest Rate Differential (IRD) = (0.0275-0.0525)

Interest Rate Differential (IRD) = -0.025

Foreign Currency Risk Premium = -0.0275 - (-0.025)

Foreign Currency Risk Premium = -0.0025 or -0.25%

b) formula, where y is the USD return and z is the percentage of Expected appreciation in foreign currency

[(1 + y) * (1 + z)] – 1

=[(1+0.0525) * (1-0.0275)]-1

=(1.0525*0.9725)-1

=1.023556-1

=0.023556 or 2.3556%